### v2

@@ 6. Threats to Validity
To put our study into place, we discuss its validity concerning the four common threat categories of construct, external, internal, and conclusion validity (Jedlitschka et al., 2008). Overall, we are confident, that our methodology allowed us to identify all security domains, where a significant amount of DSLs have been published and that we include its most influential DSLs.
Construct Validity
A potential threat to the validity of this survey is that our methodology may have missed relevant domains that do not match our search terms and are not linked to, and thus referenced by, the papers we included. However, all DSLs known to us from our research experience were covered by this systematic approach, indicating that our search terms were chosen well.
The primary focus of the survey was put onto academic papers, although organizations in practice may be the main drivers of security DSLs. Due to this decision on the survey design, some media that are often used in practice, such as blogs discussing security DSLs, and that might better reflect practice, are not covered. However, since such works are not peer-reviewed, their quality is questionable and therefore excluded with good reason. Also, such practical applications are the subject of research that is explicitly covered by the search criteria used in this work.
Using Semantic Scholar as the literature database for the search may affect the works that can be found. However, Semantic Scholar is a database that aggregates results from all major publishers. Therefore, the impact of possible threats due to the choice of Semantic Scholar is minimal.
Internal Validity
@@
There might be a bias due to the research background of the authors conducting the study, which could have impacted the execution of the study, e.g., decisions on what is considered as a security DSL. To mitigate this threat, two authors with different backgrounds, the first author with a background in cryptography and side-channel attacks and the second author in security by design, conducted each step of the investigation.
A second internal threat could arise from possible misunderstandings of the reviewed papers by the authors, which could lead to incorrect classifications. To mitigate this threat, in cases where the authors were unsure about a classification, the authors actively discussed these classifications, with other authors reading the papers in question. In addition, when analyzing the results, the authors actively read papers where the extracted statements and corresponding classifications did not seem to match. However, only minor discrepancies were identified, all of which were borderline, indicating a good quality of the extracted classifications.
External Validity
The main threat to external validity is that the papers analyzed may not be representative of security DSLs applied in practice or of all security DSLs considered in the security research community. However, the significant number of papers examined and security DSLs extracted reduces the potential impact of this threat. Furthermore, the security DSLs examined come from many different research communities and cover both scientific and practitioner-driven security DSLs, which indicates a good generality of the results.
Conclusion Validity
@@
In practice, DSLs may be more deeply integrated than observed but not yet investigated by the research community. Therefore, this could be considered the main threat to the conclusions drawn from this study. However, since the statistics of the examined security DSLs show that most of them are relatively old, but our search also covers newer works, this is unlikely.
@@ 7. Related Work
Only a few works have systematically examined security-related DSLs, and most of them have a narrower scope than this survey. Two existing works (do Nascimento et al., 2012; Kosar et al., 2016) give an overview of the domains, where DSLs have been used, but are not security specific. On the other hand, multiple surveys review tools, frameworks and DSLs for single security subdomains. Erata et al. (Erata et al., 2023) review tools for formal verification of security properties in software and hardware and include domain-specific hardware description languages for secure information flow. Tools and DSLs for the development, specification and implementation of cryptography have been evaluated in the context of computer-aided cryptography (Barbosa et al., 2021). Hastings et al. (Hastings et al., 2019) systematize the very specific cryptographic subdomain of multi party computation. Another area, that induced many new DSLs in the last years are smart contract languages and this area has been contextualized by two recent works (Varela-Vaca and Quintero, 2022; Vacca et al., 2021). A subset of the DSLs that we cover are included in publications specifically about policy specification languages (Han and Lei, 2012; Kasem-Madani and Meier, 2015; Leicht and Heisel, 2019). Also, from the domain of security-aware system modeling, DSLs have been surveyed (Hachem et al., 2016). Lallie et al. (Lallie et al., 2020) review the visual syntax of attack graphs. We identified four surveys (Uzunov et al., 2012; Nguyen et al., 2015; van Den Berghe et al., 2017; Mashkoor et al., 2023) on model-driven security engineering that mention various design-time security DSLs but mainly focus on model-based security engineering processes. A systematic overview and comparison, such as in this work, is not provided.
@@ 8. Conclusion
With this work, we present the first holistic survey of domain-specific languages for security, by analyzing 120 security DSLs based on six research questions. In particular the security research community already proposed security DSL for a wide range of security aspects. Most DSLs, address the design or implementation phases by enhancing the expression of design models, policies or code. We conclude that the integration, combination, usability, and evaluation of security-oriented DSLs still offers many opportunities for further research.

### v3

@@ 6. Threats to Validity
Our systematic methodology contributes to our confidence that we have identified all security domains where a significant number of DSLs is published, and that we have included its most influential DSLs. Still, any empirical study is prone to threats to validity (Jedlitschka et al., 2008). A potential threat is that our methodology may have missed relevant domains, especially domains that do not match our search terms and are not referenced by the papers we included. To mitigate this, we employed snowballing as recommended in ACM SIGSOFT’s Empirical Standards (Ralph, 2021). Since all DSLs known to us from our research experience, including lesser-known DSLs, were covered, this indicates that the employed methodology, search terms, and filter criteria were well chosen. We only included DSLs reported in peer-reviewed academic papers to ensure quality. However, we also considered other sources, such as project pages and wikis related to the DSLs. This ensured an informed perspective for our analysis.
@@
There might be a bias due to the research background of the authors conducting the study, which could have impacted their execution, e.g., decisions on what is considered as a security DSL. To mitigate this threat, two authors with different backgrounds conducted each step of the investigation. Still, the authors could have misunderstood the papers and done incorrect classifications. To mitigate this threat, the authors actively discussed ambiguous classifications In addition, when analyzing the results, the authors actively read papers where the extracted statements and corresponding classifications did not seem to match. Only minor discrepancies were identified and fixed.
The papers analyzed may not be representative of security DSLs applied in practice or of all DSLs considered in the security community. However, the significant number of papers and DSLs examined reduces this threat. Furthermore, they stem from many research communities, covering scientific and practitioner-driven DSLs. During our analysis, we observed that DSLs targeting the testing or deployment phases of the software development lifecycle were underrepresented in our sample. Therefore, we refrain from drawing conclusions specific to these SDLC phases and instead openly discuss possible reasons for this observation.
@@
In practice, DSLs may be more deeply integrated than observed, while not yet investigated by researchers. This could be a threat to our conclusions—but very unlikely, since our statistics show that most DSLs are relatively old, while we also covered newer works.
@@ 7. Related Work
Only few works systematically examine security-related DSLs, mostly with a narrow scope. Two works (do Nascimento et al., 2012; Kosar et al., 2016) provide an overview of the domains where DSLs have been used, but are not security-specific. On the other hand, multiple surveys review tools, frameworks, and DSLs for single security subdomains. Tools and DSLs for the development, specification, and implementation of cryptography are evaluated in the context of computer-aided cryptography (Barbosa et al., 2021). Another area that induced many new DSLs in the last years, are smart contract languages. This area is surveyed by two recent works (Varela-Vaca and Quintero, 2022; Vacca et al., 2021). A subset of the DSLs that we cover are included in publications specifically about policy specification languages (Han and Lei, 2012; Kasem-Madani and Meier, 2015; Leicht and Heisel, 2019). Also, DSLs from the domain of security-aware system modeling are surveyed (Hachem et al., 2016). Lallie et al. (Lallie et al., 2020) review the visual syntax of attack graphs. Finally, four surveys (Uzunov et al., 2012; Nguyen et al., 2015; van Den Berghe et al., 2017; Mashkoor et al., 2023) on model-driven security engineering mention various design-time security DSLs, but mainly focus on model-based security engineering processes. A systematic overview and comparison, as in our work, is not provided.
@@ 8. Conclusion
We presented the first holistic survey of 120 DSLs for security. Our focus were DSLs presented in the peer-reviewed academic literature to ensure quality and transparency. We learned that the security research community proposed security DSL for nearly every security aspect. Still, not all phases of the software development lifecycle are equally addressed. Most DSLs, address the design or implementation phases by enhancing the expression of design models, policies or code, while deployment and testing is not explicitly covered. As many DSLs specific to the underrepresented phases exist, e.g., for specifying test cases (Bisht, 2013; Peldszus et al., 2023), one reason could be that security-specific DSLs are not necessary to specify security tests. We conclude that the integration, combination, usability, and evaluation of security-oriented DSLs still offers many opportunities for further research.
Future work comprises expanding the initial framework of criteria for comparing and evaluating security DSLs. To this end, identifying reasons for the lack of security DSL adoption and derive metrics to evaluate their potential use in practice would be a valuable contribution. However, this requires a study with practitioners, which is outside the scope of this work. Finally, recall that generative AI and coding assistants have substantially enhanced software development, but do not provide guarantees about the correctness of code, let alone about the security of it (Pearce et al., 2022; Perry et al., 2023). The security DSLs we surveyed provide strong guarantees when used correctly. Finding modalities and methods of integrating the DSLs with coding assistants would be promising future research directions.